Guy Wire. How long is a guy wire reaching from the top of a 21 ft pole to a point on the ground 14 feet from the pole?
Answer by RAY100(1637)

A rough sketch shows this to be a right triangle, with legs, 14 &21, seeking the hypotenuse.
.
Pythagorous , c^2 = a^2 +b^2, applies
.
c^2 = 14^2 +21^2 = 196 + 441 =637
.
c=25.24 ft